Suppose limn→∞ (Sn − 1)/ (Sn + 1) = 0. Prove that limn→∞ Sn = 1.


1
Expert's answer
2021-02-02T04:45:07-0500

Given that "lim_{n \to \infty} \frac {S_n-1}{S_n+1}=0"

"\implies \frac {lim_{n \to \infty} (S_n-1)}{lim_{n \to \infty}( S_n+1)}=0" ........(1)

Now from (1) we can say that as the exists , therefore "lim_{n \to \infty}( S_n+1) \neq 0"

So the only possibility is that , "lim_{n \to \infty} (S_n-1)=0"

"\implies lim_{n \to \infty} S_n=1"

Which completes the proof.